OpenStudy (anonymous):

A train is on its way east was delayed 1 hr when it was 560 km west of New York City. By increasing its normal rate 10 km/hr, the train arrived at New York City on schedule. What is the train's normal rate? Solutions and answer please, I'll fan and give a medal :) thank you!

OpenStudy (anonymous):

@PaulaLovesSchool13

OpenStudy (anonymous):

@igreen @gurv

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

460km/hr i think

OpenStudy (akashdeepdeb):



OpenStudy (akashdeepdeb):

the trick is, usually, the train moves at the normal speed. let x be the normal rate or speed of the train. let d be the distance it covers before, stopping for 1 hour, 560 km before NYC. so usually, the time taken t must be, \[t = \frac{560 + d}{x}\] But in the question, after traveling properly for d kilometers, the train stops for 1 hour and then speeds up to reach at time. so, again: \[t = \frac{d}{x} + 1 + \frac{560}{x+10}\] t is the same for both, so: \[\frac{560 + d}{x} = \frac{d}{x} + 1 + \frac{560}{x+10}\] Solve this. NOTE: t is just a variable to make things simpler. d and x are required to get the answer.

OpenStudy (phi):

the hardest part (probably) is writing down equations. one way is to say: the train is traveling at x km/hr and will take t hours to go 560 km using rate*time = distance, we can write x t = 560 next, it says if we speed up (by going 10 km/h faster or x+10 ) we will "make up" an hour. that means instead of t hours it will take t-1 hours again use rate * time = distance (x+10)(t-1) = 560 both equations equal 560, so we can say x t = (x+10)(t-1) x t = xt - x +10t - 10 add -xt to both sides and we get 0 = -x +10t -10 x+10 = 10t or \[ t = \frac{x+10}{10} \] let's use that in the first equation x t = 560 \[ x \frac{x+10}{10} = 560 \] \[ x(x+10)= 5600 \\ x^2 +10x = 5600 \\ x^2 +10x -5600 = 0 \] where x is the *old speed* (or *normal* speed) i.e. what we want to find it looks like we should use the quadratic formula to solve for x.

OpenStudy (anonymous):

Thank you, I'll try both of it, and I'll give the medal to the correct one. I'll fan both of you :)

OpenStudy (akashdeepdeb):

What @phi did gives us the same equation. Both are correct.

OpenStudy (anonymous):

Yah I just finished :) anyhow, I can't give medals to both so I'll just choose later. :) Thank you so much ^_^